OpenStudy (anonymous):

Darian’s IRA has fallen 12.6% since the beginning of the year to $142,500.49. What was the value of her IRA at the beginning of the year? Please show all work

OpenStudy (anonymous):

X-0.126x=142,500.49 0.874x = 142,500.49 X=142,500.49/0.874 X=163,044.04